# A Simple and Efficient Joint Measurement Strategy for Estimating Fermionic Observables and Hamiltonians

### v1

@@ IV Gaussian unitaries for measurements of Majorana pairs and quadruples
where $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ correspond to suitably chosen permutations $R_{\pi},R_{\sigma^{(r)}}\in O(2N)$, respectively, and $U_{\mathrm{sup}}$ realizes a balanced superposition of Majorana modes belonging to disjoint subsets $Y_{\alpha}$ which form a partition of the set of all modes such that $\cup_{\alpha=1}^{L}Y_{\alpha}=[2N]$. In the rest of this section we focus on a specific case in which $2N=L(L+1)$ with the Majorana modes divided into $L$ subsets $\{Y_{\alpha}\}_{\alpha=1}^{L}$, each with cardinality $|Y_{\alpha}|=L+1$. If this relation is not satisfied, it is possible to either add $2N_{aux}\leq 2L$ auxiliary Majorana modes to embed a smaller fermionic system into a larger one, or to drop the restriction that every $Y_{\alpha}$ has the same size (see [34] for details).
The role of $U_{\mathrm{pair}}$ is to effectively realize a different pairing of the Majorana modes during the measurement process, in the sense that
@@
Figure 3: The $202N=$ 20 Majorana operators are partitioned into 5 subsets (rows) $Y^{(r)}_{\alpha}$ of cardinality 4, for each measurement setting $r=1,2$. In the first measurement setting (a), the unitary $U_{\text{sup}}$ acts via a lower flat orthogonal matrix on each row $Y_{\alpha}$, transforming every $\gamma_{k}$, $k\in Y_{\alpha}$, into a balanced superposition $\tilde{\gamma}_{k}=\sum_{l\in Y_{\alpha}}b^{(\alpha)}_{lk}\gamma_{l}$, with $b^{(\alpha)}_{lk}=\Omega(N^{-1/4})$. The pairings $(\pi(2j-1),\pi(2j))$ (marked by solid black lines) are constructed by permuting the standard pairings $\{(2j-1,2j)\}_{j=1}^{N}$ (see Fig. 1) with a permutation $\pi$ that vertically shifts (cyclically) the second and fourth columns by 1 and 2 respectively. These pairings ensure the sample from a noisy variant of $\gamma_{ij}$ can be obtained if $i$ and $j$ are in different rows. In the second setting (b), the scheme is modified by reshuffling the operators by a permutation $\sigma^{(2)}$, that shifts columns 2,3,4 (of setting 1) vertically downwards by 1,2,3 rows, respectively. All operators in the same row of setting 1 now appear in different rows of setting 2, ensuring joint measurability of all pairs.
@@
$R_{\mathrm{sup}}=\oplus_{\alpha=1}^{L}R^{(\alpha)}\ ,$ (16)
where $R^{(\alpha)}$ are operators acting on spaces $Y_{\alpha}$ constructed from $(L+1)\times(L+1)$ lower-flat orthogonal matrices defined in [34], i.e., matrices with entries satisfying $R^{(\alpha)}_{ij}=\Omega(1/\sqrt{L+1})=\Omega(1/N^{1/4})$ for $i,j\in Y_{\alpha}$. Note that orthogonal matrices with this property exist in every dimension [41].
Finally, for $r=1$ we set $U^{(r)}_{\mathrm{resh}}=\mathbb{I}$ and label $Y_{\alpha}^{(1)}:=Y_{\alpha}$. For $r>1$, the permutation $\sigma^{(r)}$ (which is not unique) acts by reshuffling the modes into different partitions $\sigma^{(r)}(Y_{\alpha}):=Y_{\alpha}^{(r)}$ of $[2N]$, for every $\alpha$.
By directly repeating steps in the proof of Prop. 1 for $U^{(r)}$ in Eq. (14), we observe that every Majorana pair $\gamma_{ij}$, with indices $i,j$ belonging to different subsets, $i\in Y_{\alpha_{i}}^{(r)},j\in Y_{\alpha_{j}}^{(r)}$, can be jointly measured with visibility $\eta_{ij}=|R^{(\alpha_{i})}_{ki}R^{(\alpha_{j})}_{lj}|$, where $(k,l)$ is the pairing between subsets $Y_{\alpha_{i}}^{(r)}$ and $Y_{\alpha_{j}}^{(r)}$ (note that by construction $l\in Y_{\alpha_{i}}^{(r)}$ and $k\in Y_{\alpha_{j}}^{(r)}$). Using the flatness property of the rotations $R^{(\alpha)}$, we get $\eta_{ij}=\Theta(1/\sqrt{N})$. By analogous reasoning, a four-element subset $A=\{ijkl\}$ together with the same unitary and pairings described above, results in $\eta_{A}=\Theta(1/N)$, as long as the individual elements $i,j,k,l$ belong to different subsets $Y_{\alpha}^{(r)}$.
@@
This modification realizes a joint measurement of all pairs and quadruples $A$ with visibilities $\eta_{ij}\geq(1/K_{2})\tilde{\eta}_{ij}$ and $\eta_{ijkl}\geq(1/K_{4})\tilde{\eta}_{ijkl}$, where $\tilde{\eta}_{A}$ is the visibility attainable from the unitary whose partition is consistent with $A$, and $K_{2},K_{4}$ are the minimal numbers of measurement settings that ensure all pairs and quadruples are compatible with at least one measurement round, respectively. In Fig. 3 we give an example of two measurement settings ($K_{2}=2$) that ensure all pairs are measured for a system of $2N=L(L+1)$ Majorana modes arranged in a rectangular layout (with $L$–even). It can also be shown that when $L+1$ is a prime number, all Majorana quadruples can be measured using $K^{\text{prime}}_{4}=7$ settings via a method that is compatible with the original layout and can be obtained by permuting Majorana modes in columns of the Majorana fermion arrangement from Fig. 3 (part a). See Appendix B for the proof. For general $L$, a probabilistic reasoning given in [34] ensures that $K_{4}=9$ random partitions of the $2N=L(L+1)$ modes into subsets of size $L$ (not necessarily compatible with the 2D layout) suffice to measure all quadruples. Furthermore, we can extend this to arbitrary $N$ mode systems by relaxing the constraint that the subsets are equally sized.
@@
The randomized version of the measurement protocol defined in Sec. III with at most 9 different FLO unitaries $U^{(1)},\ldots,U^{(9)}$ jointly measures the noisy variants $\mathbf{M}^{A,\eta_{A}}$ for all Majorana pairs and quadruples with $\eta_{A}={\Theta}(1/\sqrt{N})$ (for $|A|=2$) and $\eta_{A}={\Theta}(1/N)$ (for $|A|=4$).
@@
Importantly, in the 2D layouts from Fig. 3 (see also Fig. 4 in Sec. VI which represents a scheme for the estimation of quantum chemistry Hamiltonians) the permutations $\sigma^{(r)}$ and $\pi$ can be realized by vertical shifts of the selected columns of the arrangement of Majorana modes. This makes them feasible for experimental implementations in two-dimensional architectures in depth $\mathcal{O}(N^{1/2})$, as explained in Sec. V.
V Implementation of measurement strategies
General strategy. In this part we discuss how to implement our measurement scheme on a quantum computer. In standard fermion-to-qubit mappings (like Jordan-Wigner [39], Parity [42], Bravyi-Kitaev [37], or superfast encodings [37, 43]; see [42] for a review), unitaries from step (i) can be implemented via a depth 1 circuit consisting of Pauli gates. On the other hand, the realization of $U\in\mathrm{FLO}(N)$ from step (ii) can be achieved with a circuit of depth at most $d_{\mathrm{gen}}=\mathcal{O}(N^{3})$ and $n_{\mathrm{gen}}=\mathcal{O}(N^{4})$ two-qubit gates. This follows from two observations. First, arbitrary $U\in\mathrm{FLO}(N)$ can be realized by $U=\tilde{U}\gamma_{1}^{s}$, where $\tilde{U}$ is parity preserving (i.e., it commutes with $Q:=\gamma_{[2N]}$), and $\gamma_{1}^{s}$ ($s=0,1$) is responsible for the (optional) change of the parity. It is known that $\tilde{U}$ can be implemented using $\mathcal{O}(N^{2})$ gates of the form $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,i+1})$ [44, 45]. Second, each of the unitaries $\mathrm{exp}(\mathrm{i}\varphi\gamma_{ij})$ can be realized in depth $\mathcal{O}(N)$, which is a consequence of the fermion-to-qubit encoding that maps $\gamma_{ij}$ to a Pauli string with support on at most $\mathcal{O}(N)$ qubits, and is thus realizable in depth $\mathcal{O}(N)$ and using $\mathcal{O}(N)$ gates (unitaries $\mathrm{exp}(\mathrm{i}\varphi P)$, for a Pauli string supported on $k$ qubits can be implemented via $\mathcal{O}(k)$ two-qubit gates) [46]. However, when using the Jordan-Wigner encoding on a line it is possible to realize $U\in\mathrm{FLO}(N)$ using $\mathcal{O}(N)$ layers of $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,i+1})$ applied in parallel and forming a brickwork pattern on a triangle layout. Furthermore, in the Jordan-Wigner encoding $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,i+1})$ can be realized by a constant depth unitary acting on neighbouring qubits, and rotations corresponding to disjoint Majorana pair operators belonging to each layer can be realized in parallel (cf. [44] or [45]). This yields the overall depth $d_{\mathrm{JW}}=\mathcal{O}(N)$ and the two-qubit gate-count $n_{\mathrm{JW}}=\mathcal{O}(N^{2})$.
A square qubit grid in the Jordan-Wigner encoding. Our joint measurement scheme tailored to Majorana pairs and quadruples can be realized with a smaller circuit depth of $d_{2D}=\mathcal{O}(\sqrt{N})$ and using $n_{2D}=\mathcal{O}(N^{3/2})$ two-qubit gates in a 2D layout of qubits. Consider a $(L+1)\times L/2$ lattice of qubits realizing $L(L+1)$ Majorana modes via the Jordan-Wigner encoding (see Fig. 1, part (b)). In Sec. IV we described specific FLO unitaries $U_{\mathrm{pair}},U_{\mathrm{sup}},U^{(r)}_{\mathrm{resh}}$ that define measurement rounds of the protocol from Prop. 2. Importantly, for the variant of the Jordan-Wigner encoding in which the Majorana operators are labelled as in Fig. 3, $U_{\mathrm{sup}}$, $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ can all be implemented in depth $\mathcal{O}(N^{1/2})$ and with $\mathcal{O}(N^{3/2})$ two-qubit gates. In what follows we sketch the reasoning behind these estimates.
First, the superposition unitary $U_{\mathrm{sup}}$ corresponds to a direct sum of orthogonal transformations which independently affect the rows in which the Majorana modes are arranged in Fig. 3 (see also Eq. (16)). Such transformations are consistent with the ordering that gives rise to the JW encoding and can therefore be implemented by a (tensor) product of $L+1$ independent FLO unitaries, each acting on qubits arranged in a given row. Therefore $U_{\mathrm{sup}}$ can be implemented in depth $\mathcal{O}(N^{1/2})$ and $\mathcal{O}(N^{3/2})$ nearest-neighbour two-qubit gates.
It is more challenging to implement the unitaries $U_{\mathrm{pair}}$ and $U_{\mathrm{resh}}^{(r)}$. Note that, as explained in Section IV, permutations $\pi$ and $\sigma^{(r)}$ that correspond to unitaries $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ respectively, are “purely vertical”, in the sense that they correspond to orthogonal transformations $R_{\pi/\sigma^{(r)}}=\oplus_{\beta=1}^{L/2}\tau^{\beta}$, where $\tau^{\beta}$ only permutes elements in the $\beta$–th column. Importantly, in [44] (see also [47]), it was shown that such a class of FLO gates can be implemented in depth $\mathcal{O}(L)=\mathcal{O}(N^{1/2})$ and using $\mathcal{O}(L^{3})=\mathcal{O}(N^{3/2})$ two-qubit gates, by utilizing one extra auxiliary qubit per row (see Appendix C for the proof for general FLO transformations, the case which was not explicitly covered in [44] that focused only on number preserving Gaussian transformations).
VI Strategy for physical fermionic Hamiltonians
@@
Figure 4: Schematic representation of the construction of the 4 rounds of our measurement strategy for physical fermionic Hamiltonians from Sec. VI. The $2N$ Majorana modes are arranged on an $L\times 2L$ grid and partitioned into $L$ even ($E_{\alpha}$) and $L$ odd ($O_{\alpha}$) subsets (clusters) of cardinality $L$. The unitary $U_{\text{sup}}$ acts via a lower-flat orthogonal matrix on each cluster. In round 1, the pairings $(\pi(2j-1),\pi(2j))$ which define $U_{\text{pair}}$ (and marked by solid black lines) are constructed by a permutation $\pi$ that shifts the second and third even column cyclically by 1 and 2 respectively (in a manner analogous to Fig. 3). Importantly, the pairings (black lines) ensure each even cluster is coupled to each odd cluster. In rounds $r=2,3,4$, the different partitions of the modes into even and odd clusters are implemented by $U^{(r)}_{\text{resh}}$ via the permutations $\sigma^{(r)}$. The permutation $\sigma^{(2)}$ is realized by a vertical cyclical shift of the second and third odd columns of the first measurement setting by 1 and 2, respectively (bottom left). Similarly, $\sigma^{(3)}$ corresponds to a vertical cyclical shift of the second and third even columns of the first measurement (bottom middle). Finally, $\sigma^{(4)}$ is realized by performing both the vertical shifts realizing $\sigma^{(2)}$ and $\sigma^{(3)}$ (bottom right).
To construct an estimator for $\operatorname{tr}(H\rho)$ we propose four measurement rounds (settings). In contrast to the randomized scheme in Sec. IV, the four FLO unitaries $U^{(r)}=U^{(r)}_{\mathrm{resh}}U_{\mathrm{sup}}U_{\mathrm{pair}}$ are implemented deterministically. We consider the system size $2N=2L^{2}$ which allows for the arrangement of a rectangular $L\times L$ lattice of qubits encoding the $L^{2}$ fermionic modes (and $2L^{2}$ Majorana modes) via the Jordan-Wigner transformation (see Fig. 4 for an example for $L=3$). Systems which don’t satisfy this can always be encoded in larger ones.
The distinct aspect of this scheme is the partitioning of the $L\times 2L$ grid representing Majorana operators into even and odd subsets (clusters) labelled $E^{(r)}_{\alpha}$ and $O^{(r)}_{\alpha}$ (and coloured green and blue in Fig. 4), each containing $L$ Majorana modes. The FLO unitaries which implement the measurement are described as follows:
@@
    The permutation $\pi$ defining $U_{\text{pair}}$ translates the standard pairings to $\{(\pi(2j-1),\pi(2j))\}_{j=1}^{N}$ (marked by black lines in Fig. 4) via vertical shifts of the columns of the $L\times 2L$ grid. In particular, $\pi$ cyclically shifts the $\ell$–th even column vertically downwards by $\ell-1$ rows. The resulting pairs ensure all even clusters are connected to all odd clusters exactly once.
@@
In Appendix D we show that these four unitaries $U^{(r)}$ are sufficient to estimate all quadratic and quartic terms in $H$ with visibilities from Prop. 2. In particular, for $A\in\mathcal{X}_{2}\cup\mathcal{X}_{4}$, there exists at least one measurement round $r$ in which $A\in\mathcal{M}^{(r)}$ and therefore $\eta^{(r)}_{A}=\Theta(N^{-|A|/4})$. Note that because the FLO unitaries appearing in points (a)-(c) are either horizontal or vertical, the depth $\mathcal{O}(L)=\mathcal{O}(N^{1/2})$ (see discussion from Section IV).
@@
and $e^{(r)}_{A}\in\{\pm 1\}$ is the outcome obtained from the post-processing in step (iv) of the general protocol with unitary $U^{(r)}$ (cf. Sec. III). We impose the condition $\sum_{r}\alpha_{A}^{(r)}=1$, which ensures $\mathbb{E}\left[\sum_{r}\alpha_{A}^{(r)}\hat{\gamma}_{A}^{(r)}\right]=\mathrm{tr}(\gamma_{A}\rho)$ and therefore $\hat{H}$ is an unbiased estimator of $\mathrm{tr}(H\,\rho)$.
To improve the efficiency of our estimation strategy we can apply the median-of-means method (in analogy with classical shadows [24]). This provides a simple classical post-processing in order to reduce the effect of estimation errors. In this case the sample complexity is bounded by a multiple of the variance $\mathrm{Var}[\hat{H}]$ of the estimator, which for our scheme is given explicitly in Appendix A.
VII Numerical benchmarks

### v2

@@ II.3 Gaussian unitaries for measurements of Majorana pairs and quadruples
where $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ correspond to suitably chosen permutations $R_{\pi},R_{\sigma^{(r)}}\in O(2N)$, respectively, and $U_{\mathrm{sup}}$ realizes a balanced superposition (see Fig. 2) of Majorana modes belonging to disjoint subsets $Y_{\alpha}$ which form a partition of the set of all modes such that $\cup_{\alpha=1}^{L}Y_{\alpha}=[2N]$. In the rest of this section we focus on a specific case in which $2N=L(L+1)$ with the Majorana modes divided into $L+1$ subsets $\{Y_{\alpha}\}_{\alpha=1}^{L+1}$, each with cardinality $|Y_{\alpha}|=L$. If this relation is not satisfied, it is possible to either add $2N_{aux}\leq 2L$ auxiliary Majorana modes to embed a smaller fermionic system into a larger one, or to drop the restriction that every $Y_{\alpha}$ has the same size (see [36] for details).
The role of $U_{\mathrm{pair}}$ is to effectively realize a different pairing of the Majorana modes in measurement (1) such that
@@
$R_{\mathrm{sup}}=\oplus_{\alpha=1}^{L+1}R^{(\alpha)}\ ,$ (9)
[…]
Finally, for $r>1$, the permutation $\sigma^{(r)}$ (which is not unique) corresponding to $U^{(r)}_{\mathrm{resh}}$ acts by reshuffling the modes into different partitions $\sigma^{(r)}(Y_{\alpha}):=Y_{\alpha}^{(r)}$ of $[2N]$, for every $\alpha$. For $r=1$ we set $U^{(r)}_{\mathrm{resh}}=\mathbb{I}$ and label $Y_{\alpha}^{(1)}:=Y_{\alpha}$.
Performing the measurement with $U^{(r)}$ defined above, we can verify (by directly repeating steps in the proof of Prop. 1) that every Majorana pair $\gamma_{ij}$, with indices $i,j$ belonging to different subsets, $i\in Y_{\alpha_{i}}^{(r)},j\in Y_{\alpha_{j}}^{(r)}$, can be jointly measured with visibility $\eta_{ij}=|R^{(\alpha_{i})}_{ki}R^{(\alpha_{j})}_{lj}|$, where $(k,l)$ is the pairing between subsets $Y_{\alpha_{i}}^{(r)}$ and $Y_{\alpha_{j}}^{(r)}$ (note that by construction $l\in Y_{\alpha_{i}}^{(r)}$ and $k\in Y_{\alpha_{j}}^{(r)}$). Using the flatness property of the rotations $R^{(\alpha)}$, we get $\eta_{ij}=\Theta(1/\sqrt{N})$. By analogous reasoning, a four-element subset $A=\{ijkl\}$ together with the same unitary and pairings described above, results in $\eta_{A}=\Theta(1/N)$, as long as the individual elements $i,j,k,l$ belong to different subsets $Y_{\alpha}^{(r)}$.
@@
This modification realizes a joint measurement of all pairs and quadruples $A$ with visibilities $\eta_{ij}\geq(1/K_{2})\tilde{\eta}_{ij}$ and $\eta_{ijkl}\geq(1/K_{4})\tilde{\eta}_{ijkl}$, where $\tilde{\eta}_{A}$ is the visibility attainable from the unitary whose partition is consistent with $A$, and $K_{2},K_{4}$ are the minimal numbers of measurement settings that ensure all pairs and quadruples are compatible with at least one measurement round, respectively. In Fig. 2 we give an example of two measurement settings ($K_{2}=2$) that ensure all pairs are measured for a system of $2N=L(L+1)$ Majorana modes arranged in a rectangular layout (with $L$–even). It can also be shown that when $L+1$ is a prime number, all Majorana quadruples can be measured using $K^{\text{prime}}_{4}=7$ settings via a method that is compatible with the original layout and can be obtained by permuting Majorana modes in columns of the Majorana fermion arrangement from Fig. 2 (part a). See Appendix A for the proof. For general $L$, a probabilistic reasoning given in [36] ensures that $K_{4}=9$ random partitions of the $2N=L(L+1)$ modes into subsets of size $L$ (not necessarily compatible with the 2D layout) suffice to measure all quadruples. Furthermore, we can extend this to arbitrary $N$ mode systems by relaxing the constraint that the subsets are equally sized.
@@
The randomized version of the measurement protocol defined in Sec. II.2 with at most 9 different FLO unitaries $U^{(1)},\ldots,U^{(9)}$ jointly measures the noisy variants $\mathbf{M}^{A,\eta_{A}}$ for all Majorana pairs and quadruples with
$\eta_{A}=\begin{cases}{\Theta}(N^{-1/2})&\mbox{if}\,\,\,|A|=2\\ {\Theta}(N^{-1})&\mbox{if}\,\,\,|A|=4\,.\end{cases}$ (10)
@@
In general, in a 2D rectangular $L_{x}\times L_{y}$ layout of Majorana modes (see e.g. Fig. 2, also Fig. 3 in Sec. II.4) we can refer to transformations which act vertically or horizontally. A vertical (horizontal) FLO transformation is one whose corresponding orthogonal matrix $O$ can be written as:
$O=\bigoplus_{i=1}^{L_{x}(L_{y})}O_{i},$ (12)
where $O_{i}$ acts on modes within the $i$-th column (row) in the 2D layout. An important property of $U^{(r)}$ from (7) is that it decomposes into a product of a horizontal transformation $U_{\mathrm{sup}}$ and vertical transformations $U_{\mathrm{pair}}$, $U_{\mathrm{resh}}^{(r)}$. This makes it feasible for experimental implementations in two-dimensional qubit architectures under the JW encoding in gate depth $\mathcal{O}(N^{1/2})$ and with two-qubit gate count $\mathcal{O}(N^{3/2})$, as explained in Sec. IV.2.
II.4 Strategy for physical fermionic Hamiltonians
Figure 3: Schematic representation of the construction of the 4 rounds of our measurement strategy for physical fermionic Hamiltonians from Sec. II.4. The $2N$ Majorana modes are arranged on an $2L\times L$ grid and partitioned into $L$ even ($E_{\alpha}$) and $L$ odd ($O_{\alpha}$) subsets (clusters) of cardinality $L$. The unitary $U_{\text{sup}}$ acts via a lower-flat orthogonal matrix on each cluster. In round 1, the pairings $(\pi(2j-1),\pi(2j))$ which define $U_{\text{pair}}$ (and marked by solid black lines) are constructed by a permutation $\pi$ that shifts the second and third even column cyclically by 1 and 2 respectively (in a manner analogous to Fig. 2). Importantly, the pairings (black lines) ensure each even cluster is coupled to each odd cluster. In rounds $r=2,3,4$, the different partitions of the modes into even and odd clusters are implemented by $U^{(r)}_{\text{resh}}$ via the permutations $\sigma^{(r)}$. The permutation $\sigma^{(2)}$ is realized by a vertical cyclical shift of the second and third odd columns of the first measurement setting by 1 and 2, respectively (bottom left). Similarly, $\sigma^{(3)}$ corresponds to a vertical cyclical shift of the second and third even columns of the first measurement (bottom middle). Finally, $\sigma^{(4)}$ is realized by performing both the vertical shifts realizing $\sigma^{(2)}$ and $\sigma^{(3)}$ (bottom right).
@@ II.4 Strategy for physical fermionic Hamiltonians
To construct an estimator for $\operatorname{tr}(H\rho)$ we propose four measurement rounds (settings). In contrast to the randomized scheme in Sec. II.3, the four FLO unitaries $U^{(r)}=U^{(r)}_{\mathrm{resh}}U_{\mathrm{sup}}U_{\mathrm{pair}}$ are implemented deterministically. We consider the system size $2N=2L^{2}$ which allows for the arrangement of a rectangular $L\times L$ lattice of qubits encoding the $L^{2}$ fermionic modes (and $2L^{2}$ Majorana modes) via the Jordan-Wigner transformation (see Fig. 3 for an example for $L=3$). Systems which don’t satisfy this can always be encoded in larger ones.
The distinct aspect of this scheme is the partitioning of the $2L\times L$ grid representing Majorana operators into even and odd subsets (clusters) labelled $E^{(r)}_{\alpha}$ and $O^{(r)}_{\alpha}$ (and coloured green and blue in Fig. 3), each containing $L$ Majorana modes. The FLO unitaries which implement the measurement are described as follows:
@@
    The permutation $\pi$ defining $U_{\text{pair}}$ translates the standard pairings to $\{(\pi(2j-1),\pi(2j))\}_{j=1}^{N}$ (marked by black lines in Fig. 3) via vertical shifts of the columns of the $2L\times L$ grid. In particular, $\pi$ cyclically shifts the $\ell$–th even column vertically downwards by $\ell-1$ rows. The resulting pairs ensure all even clusters are connected to all odd clusters exactly once.
@@
In Appendix B we show that these four unitaries $U^{(r)}$ are sufficient to estimate all quadratic and quartic terms in $H$ with visibilities from Prop. 2. In particular, for $A\in\mathcal{X}_{2}\cup\mathcal{X}_{4}$, there exists at least one measurement round $r$ in which $A\in\mathcal{M}^{(r)}$ and therefore $\eta^{(r)}_{A}=\Theta(N^{-|A|/4})$. Note that because the FLO unitaries appearing in points (a)-(c) are either horizontal or vertical (see (12) for definition), they can be implemented with the same gate depth and gate count as those discussed in Sec. II.3.
@@
and $e^{(r)}_{A}\in\{\pm 1\}$ is the outcome obtained from the post-processing in step (iv) of the general protocol with unitary $U^{(r)}$ (cf. Sec. II.2). We choose coefficients $\alpha^{(r)}\in\mathbb{R}$ satisfying $\sum_{r:A\in\mathcal{M}^{(r)}}\alpha_{A}^{(r)}=1$, which ensures $\mathbb{E}\left[\sum_{r}\alpha_{A}^{(r)}\hat{\gamma}_{A}^{(r)}\right]=\mathrm{tr}(\gamma_{A}\rho)$ and therefore $\hat{H}$ is an unbiased estimator of $\mathrm{tr}(H\,\rho)$. For details of the post-processing see Appendix C. For details of the coefficient choice see Appendix E. To improve the efficiency of our estimation strategy we can apply the median-of-means method (in analogy with classical shadows [24]). This provides a simple classical post-processing in order to reduce the effect of estimation errors, and the sample complexity is bounded by a multiple of the variance $\mathrm{Var}[\hat{H}]$ of the estimator, which for our scheme is given explicitly in Appendix D.
II.5 Numerical benchmarks